OpenStudy (anonymous):

what is the slope of a line? (4,9) (2,9) a. 6 b. 9 c. unidenified d. 0

OpenStudy (anonymous):

sorry i cant put a picture, i can look for one like it maybe

pooja195 (pooja195):

Hi there! Welcome to openstudy :-) Use the slope formula \[ \huge~\rm~\bf\color{blue}{ \frac{y_2-y_1}{x_2-x_1}=slope}\]

OpenStudy (retireed):

Graph the two points to created the line, please.

OpenStudy (anonymous):

4,9 and then 2,9 is right under it

OpenStudy (anonymous):

@retirEEd

pooja195 (pooja195):

@retirEEd They want the slope the quickest way to find it is just to use the formula

pooja195 (pooja195):

\[ \huge~\rm~\bf\color{blue}{ \frac{9-9=?}{2-4=?}=slope} \]

OpenStudy (retireed):

Very good the the line is flat. Just like a table top. There is no slope by inspection.

OpenStudy (retireed):

The points with he line equation.